OpenStudy (anonymous):

Tommy, Michelle, and Allison are playing a board game. In order to determine where to move, each player has to use the spinner that has 4 colors on it. What is the probability that all three of them get red if they take turns spinning?

OpenStudy (anonymous):

required prob (as it is independent trials)=1/4 *1/4 *1/4 =1/64

Miracrown (miracrown):

i agree ^

OpenStudy (anonymous):

thank you